OpenStudy (anonymous):

each of five friends has 20 baseball cards and the decide to trade cards. molly gives jim one card, kyle gives bob one cards, molly gives kyle two cards, lisa gives molly three cards, lisa gives bob five cards, kyle gives lisa one card, molly gives bob two cards, bob gives kyle one card, and jim gives bob three cards. who has the greatest number of cards? who has the least?

OpenStudy (dumbcow):

look at net result for each person: molly: 20+(-1 -2 +3 -2) =20+(-2) = 18 kyle: 20+(-1+2-1+1) = 20+(1) = 21 jim: 20+ (1-3) = 20+(-2) = 18 bob: 20+ (1+5+2-1+3) = 20+(10) = 30 lisa: 20+(-3-5+1) =20+(-7) = 13 bob has most lisa has least

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

where did you get the negative numbers from ??

OpenStudy (dumbcow):

negative means they gave cards away positive means they received cards